# Flexible game-playing AI with AlphaViT: adapting to multiple games and board sizes

### v1

@@ 3.1 Architectures of DNNs
In AlphaViD, the move probability is estimated using the transformer decoder and MLPp. The input embeddings for the transformer decoder are generated from the outputs of the transformer encoder corresponding to the patch embeddings processed through a fully connected layer:
$E^{\prime}_{d}=MLP([\bm{z}^{2}_{L_{e}};...;\bm{z}^{n*H+m+1}_{L_{e}}]),E^{\prime}_{d}\in R^{HW\times N_{e}}.$ (7)
@@
Since the sequence size must match the input size of the transformer decoder, $E^{\prime}_{d}$ is interpolated to $E_{d}\in\mathbb{R}^{N_{a}\times N_{e}}$, where $N_{a}$ is the action space size and $N_{e}$ is the embedding size of the transformer decoder. This allows flexibility in adjusting the action size depending on the game type and board size. The transformer decoder receives $E_{d}$ and the output of the transformer encoder $z^{L}$, just like the original transformer. MLPp calculates the move probability from the output of the last layer of the transformer decoder $y_{L}$.
@@
AlphaVDA has the architecture shown on the right in Fig. 3. AlphaVDA has a similar architecture to AlphaViD, but uses learnable embeddings $E^{\prime}_{d}$ as input to the transformer decoder. These input embeddings are interpolated to fit the action size.
Figure 3: Architectures of the AlphaViD and AlphaVDA models. The input board state is divided into fixed-size patches and linearly embedded using a convolutional layer. Position embeddings are added to the resulting patch embeddings to preserve spatial information. The sequence of patch embeddings is then concatenated with a learnable value embedding and a static game embedding. This input sequence is fed into a transformer encoder. The estimated value is obtained through a multilayer perceptron (MLP) that takes the output corresponding to the value embedding from the last layer of the transformer encoder. The outputs corresponding to the patch embeddings are passed through a fully connected layer to generate input embeddings for the transformer decoder. Finally, the move probabilities are produced by another MLP that takes the output of the last layer of the transformer decoder. For AlphaVDM, the input embeddings of the transformer decoder is learnable embeddings.
3.2 AlphaViT, AlphaViD, and AlphaVDA with ResNet
AlphaViT, AlphaViD, and AlphaVDA use linear transformation to make tokens for transformer encoder. The tokenizer is not limited to linear transformation. In this study, we evaluate ResNet instead of linear transformation when it is used for tokenizer.
4 Experimental settings
@@ 4.1 Games
This study evaluates the performance of AlphaViT and AlphaViD in six games: Connect4, Connect4 5x4, Gomoku, Gomoku 6x6, Othello, and Othello 6x6. These games are two-player, deterministic, zero-sum games with perfect information. Connect4, published by Milton, is a connection game played on a $7\times 6$ board. Players take turns dropping discs onto the board. A player wins by forming a straight line of four discs horizontally, vertically, or diagonally. 54Connect4 is Connect4 with a $5\times 4$ board. Gomoku is a connection game in which players place stones on a board to form a line of five stones in a row, either horizontally, vertically, or diagonally. This study uses a $9\times 9$ board for Gomoku and a $6\times 6$ board for Gomoku 6x6. Othello (Reversi) is a two-player strategy game played on an $8\times 8$ board. In Othello, the disc is white on one side and black on the other. Players take turns placing a disc with their assigned color facing up. During a game, discs of the opponent’s color are flipped to the current player’s color if they are in a straight line and bounded by the disc just placed and another disc of the current player’s color. 66Othello is played on a $6\times 6$ board in this study.
@@ 4.2 Opponents
This study evaluates the performance of AlphaViT and AlphaViD using five different AI methods: AlphaZero, two variants of Monte Carlo Tree Search (MCTS) labeled MCTS100 and MCTS400, Minimax, and Random. AlphaZero is trained using the method described in Appendix 1. The MCTS methods (MCTS100 and MCTS400) were implemented using different numbers of simulations (100 and 400, respectively). The details of MCTS can be found in Appendix LABEL:sec:mcts. In these MCTS methods, the child nodes are expanded at the fifth visit to a node. Minimax selects a move by the minimax algorithm based on the evaluation table described in Appendix 4. Random selects moves uniformly at random from valid moves.
@@ 4.3 Software
We implemented AlphaViT, AlphaViD, the opponents, and the board games in Python, using NumPy for linear algebra and PyTorch for deep learning components. The source code is available on GitHub at https://github.com/KazuhisaFujita/AlphaViT for reproducibility and extension of this work.
@@ 5 Results
In this results, we investigate the performance and characteristics of the proposed AI game-playing agents: AlphaViT, AlphaViD, and AlphaVDA. These agents have been trained on different games (Connect4, Gomoku, and Othello) with varying board sizes (small and large). These agents employ diverse architectures in terms of the number of encoder layers to examine how the numbers of encoder layers choices impact the learning outcomes and game skill of the AI agents.
The architectures of these agents vary primarily by the number of encoder layers, which influences their capacity to learn from the game environments. The table LABEL:tab:num_params shows the number of parameters for each agent configuration. AlphaViT, AlphaViD, and AlphaVDA are tested with different numbers of encoder layers, denoted by ‘L’ followed by a number (e.g., L1, L4, L5, L8). The number of parameters ranges from 11.2 million to 19.9 million, increasing with the number of encoder layers. For comparison, the AlphaZero agent, which serves as a baseline, has 7.1 million parameters.
@@
Each AI agent is trained on specific games with varying board sizes. The table 2 categorizes the agents based on the games which they are trained on and the board sizes used during training. The first group consists of agents trained on a single game with small board size, denoted as SB. The second group includes agents trained on a single game with large board size, denoted as LB. Finally, the third group comprises agents trained on multiple games, including Connect4, Gomoku, and Othello, with large board sizes, denoted as Multi. The agents in the third group are trained simultaneously the three games. They can play these three game using only one DNN. In other words, they are not specialized for one specific game. This diversity in training regimes allows us to evaluate the agents’ adaptability and generalization capabilities across different game domains.
@@
AlphaViT SB, AlphaViD Small SB, AlphaVDA Small SB one specific game Small
AlphaViT LB, AlphaViD Small LB, AlphaVDA Small LB one specific game Large
@@
5.1 Elo rating of board game playing algorithms
@@
Table 3 shows the Elo ratings of various AI agents for different games and board sizes. Elo ratings serve as a standard measure of relative skills in two-player games and allow a systematic comparison of the performance of each AI agent. The agents include variations of the proposed methods (AlphaViT, AlphaViD, and AlphaVDM) trained on large boards (LB), small boards (SB), and multiple board sizes (Multi), as well as other AI agents, including AlphaZero, Monte Carlo Tree Search (MCTS) with different numbers of simulations, Minimax, and a Random agent. The proposed methods and AlphaZero are trained through 1000 iterations. The Elo rating is calculated through 50 round-robin tournaments between the agents. The Elo ratings of all agents are initialized to 1500. The details of the Elo rating calculation are given in Appendix 5.
@@
AlphaViT L4 LB shows strong performance, achieving Elo ratings of 1824 in Connect4, 1835 in Gomoku, and 2125 in Othello. However, for Connect4 and Gomoku AlphaViT is outperformed by AlphaZero. Interestingly, despite not being specifically trained on small boards, AlphaViT L4 LB performs as well or better than Minimax and MCTS in Connect4 5x4, Gomoku 6x6, and Othello 6x6. This suggests that AlphaViT L4 SB efficiently utilizes the knowledge obtained from large board training.
@@
AlphaViT L4 SB shows Elo ratings of 1732 in Connect4 5x4, 1764 in Gomoku 6x6 and 1969 in Othello 6x6. In these games AlphaViT L4 SB performed as well as or better than AlphaZero. For Gomoku and Othello, it outperformed Minimax and MCTS even though it is not trained on large board games. This result suggests that AlphaViT L4 SB effectively transfers knowledge from small board training to large boards.
@@
AlphaViT L4 Multi, trained simultaneously on Connect4, Gomoku and Othello, showed strong results in these games, achieving Elo ratings of 1700 in Connect4, 1969 in Gomoku and 1939 in Othello. Although it is weaker than AlphaZero in Connect4 and Gomoku, AlphaViT L4 Multi outperformed AlphaViT L4 LB (trained on Gomoku only) in Gomoku. In addition, AlphaViT L4 Multi shows comparable performance to both AlphaViT L4 SB and AlphaZero in Gomoku 6x6.
@@
AlphaViD L1 LB performs slightly worse than AlphaViT in the games with larger boards. In Gomoku 6x6 and Othello 6x6, AlphaViD L1 LB outperformed both Minimax and MCTS100, suggesting that it effectively transferred knowledge gained from training on larger boards. AlphaViD L1 SB achieved Elo ratings of 1742 in Connect4 5x4, 1785 in Gomoku 6x6 and 1852 in Othello 6x6, with performance comparable to AlphaZero and AlphaViT. AlphaViD L1 Multi showed comparable performance to AlphaViT in Connect4, but was weaker in Gomoku and Othello.
@@
Similarly, AlphaVDA L1 LB is weaker than AlphaViT for the games with large boards. AlphaVDA L1 LB outperformed Minimax and MCTS100 for all games with small board sizes. AlphaVDA L1 SB shows Elo ratings of 1770 in Connect4 5x4, 1871 in Gomoku 6x6 and 1833 in Othello 6x6, with performance comparable to AlphaZero, AlphaViT and AlphaViD. AlphaVDA L1 Multi performed better than AlphaViT in Connect4, but was outperformed by AlphaViT in both Gomoku and Othello.
@@
The Elo ratings presented in this table serve as a baseline for the subsequent experiments described in the following sections.
@@
AlphaVDA L1 SB 1213 1770 973.5 1871 1131 1833
AlphaViT L4 Multi 1700 1374 1969 1785 1939 1136
AlphaViD L1 Multi 1734 1295 1538 1654 1695 1378
AlphaVDA L1 Multi 1807 1321 1524 1194 1564 953.2
AlphaZero 2003 1767 2279 1769 1889 1769
Minimax 1033 1337 1484 1403 1127 1403
MCTS100 1260 1512 1153 1414 1192 1414
MCTS400 1550 1581 1169 1675 1365 1675
Random 725.8 1057 697.0 836.2 741.5 838.6

### v2

@@ 3.1 Architectures of DNNs
The architecture of AlphaViD’s DNN is shown at the top of Fig. 3. The DNN estimates the move probability using the transformer decoder and MLPp. The input embeddings for the transformer decoder are derived from the outputs of the transformer encoder corresponding to the patch embeddings, which are further processed through a fully connected layer as follows:
${\mathbf{y}}^{\prime}_{0}=\mathrm{MLP}([\bm{z}^{2}_{L};...;\bm{z}^{N_{e}+1}_{L}]),{\mathbf{y}}^{\prime}_{0}\in R^{N_{d}\times D_{d}},$ (7)
@@
where $N_{d}$ is the number of embeddings, and $D_{d}$ is the embedding size. Since the embedding sequence size must match the input size of the transformer decoder, ${\mathbf{y}}^{\prime}_{0}$ is interpolated to ${\mathbf{y}}_{0}\in\mathbb{R}^{N_{a}\times D}$, where $N_{a}$ is the action space size. This interpolation provides flexibility to adjust the action space size depending on the game type and board size. Similar to the original transformer, the transformer decoder receives ${\mathbf{y}}_{0}$ and the output of the transformer encoder ${\mathbf{z}}_{L}$. Finally, MLPp calculates the move probabilities from the output of the last layer of the transformer decoder ${\mathbf{y}}_{L}$. The probability of action $a_{n}$ is given by:
$p(a_{n})=\mathrm{Sigmoid}(\mathrm{MLP}_{p}(\mathrm{LN}({\mathbf{y}}_{L})))(n),$ (8)
where $a_{n}$ is the $n$th action.
@@
The architecture of AlphaVDA’s DNN is shown at the bottom of Fig. 3. The DNN has a similar architecture to AlphaViD’s DNN but uses learnable embeddings ${\mathbf{y}}^{\prime}_{0}$ as input to the transformer decoder. While the sizes of these learnable embeddings are fixed, they are interpolated to ${\mathbf{y}}_{0}$ to match the action space size, ensuring compatibility and adaptability across different game configurations.
To decide on a move, AlphaViD and AlphaVDA use Monte Carlo Tree Search (MCTS) with Upper Confidence Bound (UCT), using the value and move probabilities predicted by their DNNs. The MCTS algorithm implemented in these agents is identical to that of AlphaZero, as detailed in Appendix 1. The hyperparameters for AlphaViD and AlphaVDA are listed in Appendix 3.
4 Experimental setup
@@ 4.1 Games
This study evaluates the performance of AlphaViT and AlphaViD across six games: Connect4, Connect4 5x4, Gomoku, Gomoku 6x6, Othello, and Othello 6x6. These games are two-player, deterministic, zero-sum games with perfect information. Connect4, published by Milton, is a connection game played on a $7\times 6$ board. The players take turns dropping discs onto the board. A player wins by forming a straight line of four discs horizontally, vertically, or diagonally. Connect4 5x4 is a variant of Connect4 with a $5\times 4$ board. Gomoku is a connection game in which players place stones on a board to form a line of five stones in a row, either horizontally, vertically, or diagonally. This study uses a $9\times 9$ board for Gomoku and a $6\times 6$ board for Gomoku 6x6. Othello (Reversi) is a two-player strategy game played on an $8\times 8$ board. In Othello, the disc is white on one side and black on the other. Players take turns placing a disc with their assigned color facing up. During a game, discs of the opponent’s color are flipped to the current player’s color if they are in a straight line and bounded by the disc just placed and another disc of the current player’s color. Othello 6x6 is played on a $6\times 6$ board in this study.
@@ 4.2 Opponents
This study evaluates the performance of AlphaViT, AlphaViD, and AlphaVDA using five different AI methods: AlphaZero, two variants of Monte Carlo Tree Search (MCTS) labeled MCTS100 and MCTS400, Minimax, and Random. AlphaZero is trained using the method described in Appendix 1. The MCTS methods (MCTS100 and MCTS400) were implemented using different numbers of simulations (100 and 400, respectively). The details of MCTSs are provided in Appendix 4. In these MCTS methods, the child nodes are expanded at the fifth visit to a node. Minimax selects a move using the minimax algorithm based on the evaluation table described in Appendix 5. Random selects moves uniformly at random from valid moves.
@@ 4.3 Software
AlphaViT, AlphaViD, AlphaVDA, the opponents, and the board games were implemented in Python, using NumPy for linear algebra operations and PyTorch for deep neural networks. The source code is available on GitHub at https://github.com/KazuhisaFujita/AlphaViT for reproducibility and further extension of this work.
@@ 5 Results
The results section presents the performance and characteristics of the proposed AI agents: AlphaViT, AlphaViD, and AlphaVDA. These agents are trained on different games (Connect4, Gomoku, and Othello) with two board sizes (small and large). The architectures of these agents vary primarily according to the number of encoder layers, which influences their capability to learn from the game environment. Table 1 shows the number of parameters for each agent configuration. AlphaViT, AlphaViD, and AlphaVDA are tested with different numbers of encoder layers, denoted as ‘L’ followed by a number (e.g., L1, L4, L5, L8). The number of parameters ranges from 11.2 million to 19.9 million, increasing with the number of encoder layers. For comparison, the AlphaZero agent, which serves as the baseline, has 7.1 million parameters.
Each AI agent is trained on specific games with different board sizes. Table 2 categorizes the agents based on the games on which they are trained and the board sizes used during training. The first group includes agents trained on a single game with a large board, denoted as LB. The second group consists of agents trained on a single game with a small board, denoted as SB. The third group comprises agents trained on multiple games, including Connect4, Gomoku, and Othello, with large boards, denoted as Multi. The agents in the third group are trained simultaneously on the three games and can play these three games using a single DNN. In other words, they do not specialize in a specific game. This diversity in training regimes allows us to evaluate the adaptability and generalization capabilities of the agents across different game domains.
@@
AlphaViT LB, AlphaViD LB, AlphaVDA LB one specific game Large
AlphaViT SB, AlphaViD SB, AlphaVDA SB one specific game Small
@@
5.1 Baseline Elo ratings
@@
Table 3 lists the Elo ratings of various AI agents for different games and board sizes. Elo ratings provide a standard measure of relative performance in two-player games and enable systematic comparison of the performance of each AI agent. The evaluated agents include variations of the proposed methods trained on a single game with a large board (LB), a single game with a small board (SB), and multiple games with large boards (Multi). These are compared with other AI agents, including AlphaZero, Monte Carlo Tree Search (MCTS) with different numbers of simulations, Minimax, and a Random agent. The proposed agents and AlphaZero are trained over 1000 iterations. The Elo ratings of all agents are initialized to 1500 and calculated through 50 round-robin tournaments between the agents. The details of the Elo rating calculation are provided in Appendix 6.
@@
AlphaViT L4 LB achieves strong performance, achieving Elo ratings of 1824 in Connect4, 1835 in Gomoku, and 2125 in Othello. However, for Connect4 and Gomoku, AlphaZero outperforms AlphaViT. Interestingly, despite not being trained on small-board games, AlphaViT L4 LB performs as well as or better than Minimax and MCTS100 in Connect4 5x4, Gomoku 6x6, and Othello 6x6. Furthermore, AlphaViT L4 LB exhibits competitive performance for Othello with AlphaViD L1 SB, AlphaVDA L1 SB, and AlphaZero. This suggests that AlphaViT may efficiently utilize the knowledge obtained from training on large-board games.
@@
AlphaViT L4 SB shows Elo ratings of 1732 in Connect4 5x4, 1764 in Gomoku 6x6, and 1969 in Othello 6x6. In these games, AlphaViT L4 SB performs as well as or better than AlphaZero. For Gomoku and Othello, it outperforms Minimax and MCTS even though it is not trained on large board games. This result suggests that AlphaViT L4 SB may effectively utilize knowledge from training on small-board games for large-board games.
@@
AlphaViT L4 Multi, trained simultaneously on Connect4, Gomoku, and Othello, shows strong results in these games, achieving Elo ratings of 2003 in Connect4, 1969 in Gomoku, and 1939 in Othello. Although it is weaker than AlphaZero in Connect4 and Gomoku, AlphaViT L4 Multi outperforms AlphaViT L4 LB (trained on Gomoku only) in Gomoku. Furthermore, AlphaViT L4 Multi exhibits comparable performance to MCTS400 in Gomoku 6x6.
@@
AlphaViD L1 LB performs slightly worse than AlphaViT in the games with the large boards. For Gomoku 6x6 and Othello 6x6, AlphaViD L1 LB outperforms Minimax and MCTS100, suggesting that it effectively transfers knowledge gained from training on large boards. AlphaViD L1 SB achieves Elo ratings of 1742 in Connect4 5x4, 1785 in Gomoku 6x6, and 1852 in Othello 6x6, with performance comparable to AlphaZero and AlphaViT. AlphaViD L1 Multi shows performance close to AlphaViT in Connect4 but is weaker in Gomoku and Othello.
@@
AlphaVDA L1 LB performs worse than AlphaViT for the games with large boards. AlphaVDA L1 SB shows Elo ratings of 1770 in Connect4 5x4, 1871 in Gomoku 6x6, and 1833 in Othello 6x6, with performance comparable to AlphaZero, AlphaViT, and AlphaViD. AlphaVDA L1 Multi is on par with or slightly weaker than AlphaViD L1 Multi.
@@
The Elo ratings presented in Tab. 3 serve as a baseline for the subsequent experiments described in the following sections.
@@
AlphaVDA L1 SB 1213 1770 973.5 1785 1131 1833
AlphaViT L4 Multi 2003 1374 1969 1654 1939 1136
AlphaViD L1 Multi 1743 1295 1538 1392 1695 1379
AlphaVDA L1 Multi 1720 1321 1524 1194 1565 953.2
AlphaZero 2058 1767 2279 1769 1889 1850
Minimax 1033 1337 1484 1403 1127 1385
MCTS100 1260 1512 1153 1414 1192 1372
MCTS400 1550 1581 1169 1675 1365 1642
Random 725.8 1057 697.0 838.6 741.5 883.2
5.2 Variation of Elo rating over training iteration
Figure 4 shows the Elo ratings of the AlphaViT variants (AlphaViT, AlphaViD, AlphaVDA) and AlphaZero over multiple iterations for large and small board configurations in three games: Connect4, Gomoku, and Othello. The left column shows the results for the large board configurations (standard sizes for Connect4, Gomoku, and Othello). Here, the Elo ratings for single-game-trained agents (AlphaViT L4 LB, AlphaViD L1 LB, AlphaVDA L1 LB), multi-game-trained agents (AlphaViT L4 Multi, AlphaViD L1 Multi, AlphaVDA L1 Multi), and AlphaZero were evaluated at iterations 1, 2, 4, 6, 8, 10, 20, 40, 60, 80, 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000, 1100, 1200, 1300, 1400, 1500, 1600, 1700, 1800, 1900, 2000, 2100, 2200, 2300, 2400, 2500, 2600, 2700, 2800, 2900, and 3000. The right column displays the performance on small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6), with evaluations of single-game-trained models (AlphaViT L4 SB, AlphaViD L1 SB, AlphaVDA L1 SB) and AlphaZero at iterations 1, 2, 4, 6, 8, 10, 20, 40, 60, 80, 100, 200, 300, 400, 500, 600, 700, 800, 900, and 1000.
For games with large boards, all single-game-trained models (AlphaViT L4 LB, AlphaViD L1 LB, AlphaVDA L1 LB) show rapid improvements in Elo ratings, reaching high levels around 300 iterations for Othello and 500 iterations for Gomoku. For Connect4, the Elo ratings continue to improve steadily, with gains even after 1000 iterations. The multi-game-trained models (AlphaViT L4 Multi, AlphaViD L1 Multi, and AlphaVDA L1 Multi) exhibit a similar trend in their Elo ratings to the single-game-trained models. Their performance across the different games shows the flexibility of these agents. After the initial rapid improvement phase, the agents show only minor gains, and their Elo ratings fluctuate.
For small boards, AlphaViT L4 SB, AlphaViD L1 SB, and AlphaVDA L1 SB show very rapid improvements in Elo ratings. After peaking, they reach stable Elo ratings for Connect4 5x4 and Gomoku 6x6. However, for Othello 6x6, the Elo ratings exhibit more fluctuations even after reaching high-performance levels. Because the Elo ratings of the agents rapidly increase and become stable, the evaluation iterations are set to be fewer than those used for larger boards.
These results show that AlphaViT, AlphaViD, and AlphaVDA perform well on small boards, achieving high and stable Elo ratings. However, for games with large boards, agents’ Elo ratings exhibit larger fluctuations. This suggests that further optimizations, such as increasing the depth of the DNNs or adjusting hyperparameters, may be necessary to enhance game-playing skills on larger boards.
Figure 4: Elo rating progression over training iterations for AlphaViT, AlphaViD, and AlphaVDA in large- and small-board games. The left column shows results for large board configurations (Connect4, Gomoku, and Othello), while the right column displays results for their small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6). Solid lines represent single-game-trained agents, whereas dashed lines indicate multi-game-trained agents.
5.3 Effect of transformer encoder depth on performance
In this subsection, we evaluate the performance of agents with deep encoders: AlphaViT L8, AlphaViD L5, and AlphaVDA L5. Figure 5 shows the variations in Elo ratings for these agents over the iterations. The opponents and evaluation conditions are consistent with those used in previous experiments. For the large board configurations (Connect4, Gomoku, and Othello), the Elo ratings stabilize at about 1000-2000 iterations. In contrast, the Elo ratings stabilize at around 100 iterations for Connect4 5x4 and 200 iterations for Gomoku 6x6. For Othello 6x6, the Elo ratings quickly reach a high level but exhibit some fluctuations.
Table 4 provides the mean Elo ratings for each agent, calculated from 2100 to 3000 iterations for the large board configurations. Figure 6 displays the ratio of the mean Elo rating of the agents with deep encoders to that with shallow encoders. The results demonstrate that deeper architectures generally enhance the performance across all games. Notably, AlphaViD Multi and AlphaVDA Multi show significant improvements, particularly for Gomoku, which reach ratios of 1.220 and 1.165, respectively. These agents also improve the performance for Connect4, with ratios of 1.130 and 1.143, and for Othello, with ratios of 1.105 and 1.146. The single-game-trained AlphaViD and AlphaVDA exhibit significant gains for Connect4 and Othello, while their improvements for Gomoku are more modest. In contrast, the single- and multi-game-trained AlphaViTs show weaker improvements than the other agents. These results suggest that increasing the depth of transformer encoder layers enhances the agent’s performance across different board configurations and game types, with a particularly positive impact on AlphaViD and AlphaVDA.
Figure 5: Elo rating progression over training iterations for AlphaViT, AlphaViD, and AlphaVDA with deep encoder in large- and small-board games. The left column shows results for large board configurations (Connect4, Gomoku, and Othello), while the right column displays results for their small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6). Solid lines represent single-game-trained agents, whereas dashed lines indicate multi-game-trained agents.
Table 4: Mean Elo Ratings from 2100 to 3000 iterations
Game Connect4 Gomoku Othello
AlphaViT L4 LB 1966 1753 2007
AlphaViT L8 LB 2028 1862 2046
AlphaViD L1 LB 1814 1656 1756
AlphaViD L5 LB 1902 1704 1928
AlphaVDM L1 LB 1802 1570 1807
AlphaVDM L5 LB 1996 1602 1992
AlphaViT L4 Multi 1897 1917 1964
AlphaViT L8 Multi 2015 2037 1965
AlphaViD L1 Multi 1690 1611 1711
AlphaViD L5 Multi 1909 1966 1891
AlphaVDA L1 Multi 1721 1583 1686
AlphaVDA L5 Multi 1968 1844 1933
AlphaZero 2068 2187 1890
Figure 6: The ratio of Elo ratings between deep and shallow DNNs for AlphaViT, AlphaViD, and AlphaVDM across Connect4, Gomoku, and Othello. Ratios greater than 1.0 indicate superior performance of agents with deep DNNs compared to those with shallow DNNs.
5.4 Effect of fine-tuning from small board games
Figure 7 illustrates the Elo ratings of AlphaViT L8, AlphaViD L5, and AlphaVDA L5 with fine-tuned DNNs and those without fine-tuning (initialized with random weights) across three games: Connect4, Gomoku, and Othello. The fine-tuned DNNs are initialized using weights from the single-game-trained DNNs trained on the small-board configuration at 200 iterations. For example, the fine-tuned DNN for Connect4 is initialized with weights from the DNN trained on Connect4 5x4. In contrast, agents with non-fine-tuned DNNs are initialized with random weights.
The results demonstrate that the agents with the fine-tuned DNNs consistently achieve higher Elo ratings than those with randomly initialized DNNs, particularly for Gomoku and Othello. The performance difference is pronounced in the early training iterations, with fine-tuned DNNs exhibiting rapid improvements and stabilization for Gomoku and Othello. For Gomoku, the advantage of fine-tuning is obvious, as the agents with the fine-tuned DNNs consistently outperform those with the non-fine-tuned DNNs throughout the training process. Some signs of overfitting appear in AlphaViD L5 and AlphaVDA L5 for Gomoku, as their Elo ratings gradually decline from approximately 1000 to 2000 iterations. However, their performance improves after 2000 iterations. In Connect4, the impact of fine-tuning is less significant, with no substantial difference in Elo ratings between agents using fine-tuned and non-fine-tuned DNNs.
Figure 7: Elo rating progression over training iterations of the agents (AlphaViT L8, AlphaViD L5, and AlphaVDA L5) with finetuned and randomly initialized DNNs for three board games (Connect4, Gomoku, and Othello). The three columns show the results for AlphaViT L8, AlphaViD L5, and AlphaVDA L5, respectively. The three rows show the results for Connect4, Gomoku, and Othello.